The following is a list of account titles and amounts (in millions) reported at December 27, 2015, by Hashey, Inc a leading manufacturer of games, toys, and interactive entertainment software for children and families: $1,096 Equipment s 476 Accumulated Amortization Accumulated Depreciation Allowance for Doubtful Accounts Buildingr 731 Goodwill 486 Inventortes 40 Land 226 Licensing Rights 666 Prepaid Rent 336 Requlred 1. Prepare the asset section of a classified balance sheet for Hashey Inc. 2. Using Hasheys 2015 Net Sales Revenue of $4620 (million) and its Net Fixed Assets of $216 (million) at December 28 2014 calculate the fixed asset turnover ratio for 2015. Has the company generated more or less revenue from each dollar of fixed assets than in 2014. when the ratio was 20 423
Answer #1

Answer: Hashey, Inc. Balance Sheet (Partial) December 27, 2015 (in Millions ASSETS Current Assets: Cash and Cash Equivalents Accounts Receivable Less: Allowance for Doubtful Accounts Inventories Prepaid Rent Total Current Assets Property, Plant and Equipment: Equipment Buildings Land Property, Plant and Equipment (At Cost) Less: Accumulated Depreciation Property, Plant and Equipment, Net Intangible Assets: Goodwill Licensing Rights Intangible Assets (At Cost) Less: Accumulated Amortization Intangible Assets, Net Total Assets 1096 1056 40336 341 2399 476 226 486 225 581 1869 -2450 731 1719 4343

Answer to Part 2.

Fixed Assets Turnover = Sales / Average Fixed Assets

Average Fixed Assets = (216 + 225)/2 = $220.50 million

Fixed Assets Turnover = 4,620/ 220.50

Fixed Assets Turnover = 20.95

The company has generated more of Revenue.